# Latent Policy Steering with Embodiment-Agnostic Pretrained World Models

### v2

@@ IV-B Latent Policy Steering with a world model
11:Optimize the value function $\mathcal{V}$ with the loss: $\mathcal{L}(R_{\text{all}},V(S_{\text{all}}))$.
Inference with LPS
A large batch of action plans is sampled from the BC policy during inference, as shown in phase 3 of Fig. 2. We use a batch size of $b=400$. After encoding the current observation, the WM predicts the corresponding future states for each plan and computes the state values. A value is assigned to each plan based on a weighted average of state values with a discount factor, by assigning heavier weights to future states. The motivation is to accommodate the fact that the value function is not perfect. In practice, we found that this discounted weighted averaging of values works better than using the average or the last-state value per plan to represent the plan-level value. The action plan with the highest value is executed in open loop.
IV-C Algorithm summary
Our approach includes 3 phases: 1) We curate a pre-training video dataset including different robots and humans completing manipulation tasks. We compute optical flow between frames as actions and pretrain a WM with optical flow as the action representation. 2) We collect a small robot dataset on the robot to learn a policy from scratch. We finetune the WM on the robot dataset with robot action (discarding the optical flow encoder). We also learn a robust value function for LPS (Alg. 1). Lastly, 3) We improve the action selection during inference by policy steering with the robust value function. An illustration is provided in Fig. 2.
@@ V Experiments
The proposed method is evaluated on object manipulation tasks from a simulated benchmark (Robomimic [37]) and real world environments. We compared our approach to baselines with a pretrained policy and learning from scratch. Ablations are provided to verify the key choices of the algorithm and the effectiveness of the embodiment-agnostic actions across a wide range of embodiments. In the following, we first describe our experiment settings, including environment, datasets, and baselines.
@@ V-A Settings
Environment: Robomimic. Our approach and baselines are evaluated on a Franka robot across 3 tasks (Lift, Can, Square) from Robomimic [37]. Instead of leveraging 300 expert demonstrations provided by the benchmark, we only use the 30-50 expert demonstrations to learn/finetune our models, mimicking a low-data regime setting. For model pretraining, we use Robosuite [38] with a joystick to collect 100 demonstrations on 3 robots other than Franka: IIWA, UR5e, and Kinova3 via the Robosuite interface. In this case, the pretraining data has action labels inherently different from those of the Franka robot. Our evaluation protocol is similar to [39]. We report the average success rates and standard deviations across 3 seeds. Each seed has a success rate based on hundreds of evaluations.
@@
Environment: Real world. We create an evaluation environment with a Franka robot, a desk, and objects to manipulate (cups, pot, spoon, oven, towel, beads, pan, toy radish), shown in Fig. 4. We mount a wrist camera on the robot’s wrist and a fixed camera on the side. For each task, only the objects involved in the task will be presented. The target dataset is collected via teleoperation with joystick. The positions of the objects during data collection are randomly selected from the desk area. The object positions during evaluations include 20 positions that are pre-selected, uniformly covering in the desk area. We consider five tasks, with three basic manipulation tasks (put-radish-in-pot, stack-cups, open-oven-put-pot-in-close-oven), and two challenging long-horizon tasks that require tool use or precision (scoop-beads-with-spoon, fold-towel-to-triangle). Scooping requires the robot to use a spoon to move small beads from a pan to a bowl. Imprecise manipulation will easily lead to beads falling out of the spoon. Folding requires multiple pick-and-place actions to align the towel’s corners.
Figure 4: Real world environment. We conduct the real-world experiment in a table-top setting with a Franka robot, given a set of common objects.
@@
Datasets. We use pretraining datasets and target datasets in our experiment. The target dataset is a small Franka dataset with robot actions (e.g., end-effector pose) and expert quality (all demonstrations succeed and result in task completions), used to learn a policy and finetune the WM. The pretrain dataset is used to pretrain the WM, without assuming expert quality. It was made of data collected in simulation, existing public robot datasets, or easy-to-collect human data. In our case, we ask the human to interact with the desk environment without a specific purpose, known as data from play. We consider 3 pretraining data variants:
@@
  • -sim: A dataset of simulation robot data collected by a human teleoperator. It includes 3 robots, different from the robot used for evaluation (target robot) with 100 demonstrations for each robot.
@@
  • -real: A set of public robot data collected by different research groups in the real world ( Open X-Embodiment [4]). Since most of the object manipulation data comes from several popular robots in the Open X-embodiment dataset, we filter data from 5 popular robots. The resulting dataset has 2000 episodes after dropping some short or long episodes.
@@
  • -human: A dataset made of human video data from play. A human is asked to interact with the environment without specific goals. Therefore, frequent unintentional task completions/resets throughout the playing episode contain richer information to learn, compared to an episode collected via teleoperations. It’s total transitions is about 43% of the public robot dataset.
@@
  • BC: a diffusion policy [2] without pretraining. It is learned via behavior cloning on the target dataset. All of our experiments use an action-chunking size of 16.
  • HPT: a pretrained policy [22] finetuned on the target dataset. The transformer backbone is frozen, and the action prediction head and the low-level encoder are finetuned. It uses an action chunking size of 16.
  • LPS*: our proposed method (LPS) without pretraining the WM. The WM is trained only on the target dataset. The WM will be combined with the policy (BC), and has a prediction horizon of 16. The best action plan, according to the WM, is executed in open loop.
  • LPS-(sim/real/human/mix)*: LPS with WMs pretrained on robot data from simulations (-sim), public robot data collected in the real world (-real), human videos collected in the real world(-human), or the mixture of -sim/real/human. It has a prediction horizon of 16. The best action plan, according to the WM, is executed in open-loop
@@
V-B Comparisons in the simulation & real world
We compared LPS’s variants without pretraining (LPS*) and with pretraining (LPS-sim/real/human/mix*) against a behavior cloned policy (BC) across tasks both in the Robomimic and real world. LPS uses the exact same policy as the BC baseline. The main difference is that LPS uses a WM and a robust value function to select the best action from the BC.
Figure 5: Effects of the pretrain embodiments. World Models (WM) are pretrained on data with different embodiments (i.e., recipes). They are finetuned on the same robot dataset and combined with the same policy to solve the same task. We observe a promising trend: our approach can effectively leverage data across multiple embodiments. By increasing the number of pretraining embodiments, the amount of pretraining data grows, results in better performance compared to the policy-only baseline (BC) and WM without pretraining (i.e., LPS-scratch). Surprisingly, human video data leads to a competitive WM checkpoint (green), compared to the recipe with more embodiments and more data.
In general, the LPS with a pretrained WM (LPS-sim/real/human/mix*) consistently improves the policy (BC) performance across tasks both in the simulation and real world (Table I, Table II). Thanks to the embodiment-agnostic action representation via optical flow, increasing the embodiments in the dataset results in a larger usable dataset. This advantage translates into better overall performance.
TABLE I: Latent Policy Steering with WM in Robomimic
@@ V-B Comparisons in the simulation & real world
Average 36.6% 37.8% 40.0% 39.1% 39.6% 67.8% 60.0% 74.9% 72.9% 73.1%
We report the mean success rate across 3 seeds with standard deviations. The proposed method (LPS, LPS-sim*/real*/mix*) can greatly improve
a policy (BC) via a WM during inference, in low-data regime. We found not only the simulation data helps to improve the performance (LPS-
sim*) via WM pretraining, real-world robot and human data improves the WM via pretraining as well (LPS-real*/mix*).
TABLE II: Latent Policy Steering with WM in the Real World
[…]
BC LPS* HPT
LPS
-real*
LPS
-human*
LPS
-mix*
BC LPS* HPT
LPS
-real*
LPS
-human*
LPS
-mix*
Put-radish-
in-pot
14/20a 15/20a 5/20a 16/20a 17/20a \ 16/20c 18/20c 10/20c 20/20c 20/20c \
Stack-cups 10/20a 14/20a 2/20a 14/20a 16/20a \ 14/20c 14/20c 3/20c 15/20c 17/20c \
Open-oven-
put-pot-in-
close-oven
2/20a 6/20a 0/20a 9/20a 8/20a \ 8/20c 11/20c 0/20c 12/20c 11/20c \
Average
43.3% 58.3% 11.7% 65.0% 68.3% \ 63.3% 71.7% 21.7% 78.3% 80.0% \
Scoop-beads-
with-spoon
6/20b 5/20b 0/20b \ \ 10/20b 13/20d 16/20d 0/20d \ \ 16/20d
Fold-towel-
to-triangle
0/20b 0/20b 0/20b \ \ 2/20b 9/20d 11/20d 3/20d \ \ 13/20d
Average
15.0% 12.5% 0.0% \ \ 30.0% 55.0% 67.5% 7.5% \ \ 72.5%
We report the number of successes out of 20 trials in the real world. Given challenging tasks with limited data, our method improves a policy
(BC) during inference via a WM (LPS*, LPS-real*/human*/mix*). Specifically, pretraining the WM on different embodiments (LPS-real*/hum-
an*/mix*) leads to better results than the WM learned from scratch (LPS*).
V-C Comparisons of different embodiments for pretraining
We explore how different combinations/numbers of embodiments for pretraining affect the performance during evaluations on the Can task from Robomimic. Our target dataset has 50 demonstrations, and the pretraining dataset could include simulation-only robot data, real-world robot data, and human video data from play. The BC baseline is a behavior cloning policy, and LPS-scratch uses a WM learned from scratch to improve the BC policy. Other LPS variants use different recipes of embodiments during pretraining. We observe a promising trend: due to our embodiment-agnostic action representation that allows us to leverage cross-embodiment data, increasing the number of embodiments leads to more usable data and results in better performance compared to the policy-only baseline (bc) and WM without pretraining.
Surprisingly, we observe that the human video data from play leads to a very competitive pretrained WM, despite having less data compared to other data mixtures with more embodiments. We suspect that this benefit comes from the human demonstrator, who can demonstrate more diverse and coherent manipulation skills in a shorter time with their hands compared to a teleoperation device. The resulting data are of higher quality compared to the teleoperation data, which are used to collect most of the large robot datasets.
Figure 6: Latent Policy Steering with different horizons. We found the proposed LPS method can effectively improve policy performance from a shorter horizon to a longer horizon up to 16. When the horizon becomes very long (e.g., 20, 24), the reward based on state similarity becomes noisy, making the value function less useful during inference.
@@ V-C Comparisons of different embodiments for pretraining
We compare the WM pretrained with embodiment-agnostic actions (optical flow) against the WM pretrained with end-effector action (eef). On a dataset of 3 robots, the optical-flow-based WM leads to better results. We believe the performance gap over the eef action will grow when more embodiments are provided.
V-D Comparing optical flow to EEF actions
@@
Robot end-effector pose (EEF) is another potentially embodiment-agnostic action representation for manipulators. In simulation, we investigate the advantage of using optical flow as an action representation during pretraining over EEF pose, given a pretraining dataset with 3 different robots. Our approach (LPS-sim(flow)*) uses optical flow action during pretraining, and LPS-sim(eef) uses EEF action during pretraining. Both are finetuned on the target dataset with 30 or 50 demonstrations, using the EEF robot action.
@@
The results are shown in Table III. The advantage of optical flow as embodiment-agnostic actions vs. EEF actions is clear, given 3 embodiments. We suspect the advantage will keep growing when the dataset size is larger and the number of embodiments is larger. Notice that optical flow also allows human embodiment to be part of pretraining, while EEF actions are only available for robots
V-E LPS’s performance across horizons
We investigate LPS’s performance with different action chunking horizons and the WM’s prediction horizon from small to large. The WM’s prediction horizon always matches the action chunking size. We learn the policy and the WM from scratch, given EEF action, with 100 demonstrations on the Can task of Robomimic [37].
We found that LPS with a WM performs better than BC at horizons 4,8,16,20, and performs worse than BC for a horizon of 24 (Fig. 6). Large action chunking size/horizon of 24 leads to suboptimal policy with poor performance, and the penalization on deviation from the dataset will be noisy (lines 8-10, Alg. 1). This results in a noisy reward and makes the value function less useful.
V-F Ablations on the LPS’s robust value function
LPS uses a robust value function that has seen data that are both in-distribution and out-of-distribution, and it penalizes all behaviors deviating from the dataset, where the policy may not be recovered from. We verify these ideas via 2 variants of LPS.
  • LPS-vanilla: This variant does not let the value function see the out-of-distribution data. Policy Steering with WM and vanilla value function (PS-vanilla) is a variant that learns a vanilla value function on the dataset, given environmental rewards. It will select the best candidate based on the value function.
  • LPS-bootstrap: Policy Steering with bootstrapping is a variant that learns a value function both on the dataset (in-distribution), and outside of the dataset (out-of-distribution). It achieves this by simulating the states likely visited by the policy and learns a value function. Different from LPS, the extra states are labeled with the same rewards as the dataset’s states and do not include the term penalizing deviation from the expert states ($r^{\prime}_{t:t+h}=r_{t:t+h}$ at line 8, 1). It will select the best candidate based on the value function.
All the models use 100 demonstrations to train from scratch, and no pretraining are involved.
TABLE IV: Latent Policy Steering with ablations on the value function
@@ V-F Ablations on the LPS’s robust value function
Transport 30.4±32.9 20.4±1.4 22.0±2.1 34.6±21.6
Average 63.7% 62.6% 62.3% 68.5%
We ablate our designs of the robust value function for LPS into 2 variants (LPS-vanilla/bootstrap). In both variants, performance is lower than the policy-only baseline (BC), demonstrating the necessity of designing the robust value function to let WM benefit the policy (BC).
VI Conclusion
We pretrain an embodiment-agnostic World Model (WM) on existing or cost-effective data sources such as public multi-embodiment robot datasets [4], robot data from simulations [38], and easily collected human data from play to improve visuomotor policies with a small amount of real-world data. Thanks to the optic flow as an embodiment-agnostic action representation, the pretrained WM has minimal dependency on specific embodiments and can be easily finetuned to a target embodiment for deployment with a small amount of data. Through 5 real-world tasks, our proposed method improves a policy’s relative performance during inference over 50% with 30-50 expert demonstrations, and over 25% relative improvement with expert 50-100 demonstrations, showing its effectiveness over prior work without pretraining a WM. The code for pretraining and finetuning the WM for LPS will be made publicly available.

### v3

@@ IV-B Target Embodiment Finetuning
[…]
11:Optimize the value function: $\mathcal{L}(R_{\text{all}},\mathcal{V}(S_{\text{all}}))$.
12:Updates $\mathcal{V}_{target}$ given $\mathcal{V}$.
IV-C Latent Policy Steering
We refer to our method of finetuning a WM and training a robust value function steering a policy back to the expert dataset via WM-generated latent states as Latent Policy Steering (LPS). LPS improves the base behavior cloning policy during inference by evaluating multiple candidates in the WM’s latent state space and selecting the best one according to the value function. We sample $B$ number of action plans with horizon $h$. After encoding the current observation, the WM predicts the corresponding future states for each plan and computes the state values. A plan-level value is computed based on a weighted average of state values with a discount factor, by assigning heavier weights to future states to reduce noise in the value function predictions. In practice, we found that this discounted weighted averaging of values works better than using the average or the last-state value per plan to represent the plan-level value. Finally, the action plan with the highest value is executed for the given horizon before replanning.
Figure 2: Real-world experiment. The proposed method and baselines are evaluated on: (a) Put-radish-in-pot, (b) Sweep-salad-off-the-board, (c) Scoop-beads-with-spoon, and (d) Fold-towel-to-triangle.
@@ V Experiments
We evaluate the proposed method on real-world environments and on a simulated benchmark (Robomimic [22]). We compared our embodiment-agnostic approach to baselines with an embodiment-dependent pretrained policy [35] and learning from scratch. Ablations are provided to verify the key choices of the algorithm and the effectiveness of the embodiment-agnostic actions across multiple embodiments. In the following, we first describe our experiment settings, including environment, datasets, and baselines.
@@ V-A Settings
Environment: Real world. We create an evaluation environment with a Franka robot, a desk, and objects to manipulate. We mount a camera on the robot’s wrist and a fixed camera on the side. The target dataset is collected via teleoperation with a joystick. The positions of the objects during data collection are randomly selected from the desk area. The object positions during evaluations include 20 positions that are pre-selected, uniformly covering the desk area. We consider a pick-and-place task and three challenging long-horizon tasks that require tool use or precise manipulations of deformable objects (sweep-salad-off-the-board, scoop-beads-with-spoon, fold-towel-to-triangle), shown in Fig. 2. Scooping requires the robot to use a spoon to move small beads from a pan to a bowl. Imprecise manipulation will easily lead to beads falling out of the spoon. Folding/Sweeping both require interactions to manipulate deformable objects.
Environment: Robomimic. Our approach and baselines are evaluated on a Franka robot across 4 tasks (Lift, Can, Square, Transport) from Robomimic [22]. We only use the 30-50 expert demonstrations to learn/finetune our models, mimicking a low-data regime setting. We use Robosuite [42] with a joystick to collect 100 demonstrations on 3 robots other than Franka: IIWA, UR5e, and Kinova3 via the provided interface. In this case, the pretraining data has action labels inherently different from those of the Franka robot. Our ablation study for LPS’s design choices and different horizons uses 100 demonstrations. Our evaluation protocol is similar to [29]. We report the average success rates and standard deviations across 3 seeds. Each seed has a success rate based on hundreds of evaluations.
@@
Datasets. We use pretraining datasets $\mathcal{C}$ and target datasets $\mathcal{E}$ in our experiment. The target dataset is a small expert quality (all demonstrations succeed and result in task completions) Franka dataset with robot actions (e.g., end-effector pose), which is used to learn a policy and finetune the WM. The pretrain dataset $\mathcal{C}$ has multiple embodiments, with no assumptions on quality. The pretraining dataset is composed of data collected in simulation, existing public robot datasets, or easy-to-collect human data. In our case, we ask the human to interact with the real-world environment without a specific purpose, known as data from play. We consider 3 pretraining data variants:
@@
  • -sim: A dataset of simulation robot data collected by a human teleoperator. It includes 3 robots, different from the robot used for evaluation (target robot), with 100 demonstrations for each robot.
@@
  • -real: A set of public robot data collected by different research groups in the real world ( Open X-Embodiment [34]). Since most of the object manipulation data comes from several popular robots in the Open X-embodiment dataset, we filter data from 4 popular robots. The resulting dataset has 2000 episodes after dropping episodes that are too short or too long.
@@
  • -human: A dataset made of human video data from play. A human is asked to interact with the environment without specific goals. Therefore, frequent unintentional task completions/resets throughout the playing episode contain richer information to learn, compared to an episode collected via teleoperations. Its total transitions are about 43% of the public robot dataset.
@@
  • BC: a diffusion policy [8] learned from scratch via behavior cloning on the target dataset. All of our experiments use an action prediction horizon of 16.
  • HPT: a cross-embodiment pretrained policy [35] finetuned on the target dataset, including embodiment-specific action head and the low-level encoder. It uses an action prediction horizon of 16.
  • LPS*: our proposed method without pretraining the WM. The WM is trained only on the target dataset. The WM will be combined with the policy (BC), and has a prediction horizon of 16. The best action plan, according to the WM, is executed in an open loop.
  • LPS-(sim/real/human/mix)*: LPS with WMs pretrained on robot data from simulations (-sim), public robot data collected in the real world (-real), human videos collected in the real world(-human), or the mixture of -sim/real/human. It has a prediction horizon of 16.
@@
V-B Real-world experiments
TABLE I: Latent Policy Steering in the real world
[…]
BC LPS* HPT
LPS-mix*
BC LPS* HPT
LPS-mix*
Put-radish-
in-pot
[…]
Sweep-salad-
off-the-board
[…]
Scoop-beads-
with-spoon
[…]
Fold-towel-
to-triangle
[…]
Average
21.2% 21.2% 5.0% 36.2% 51.2% 62.5% 13.8% 73.8%
We report the number of successes out of 20 trials in the real world. Thanks to the embodiment-agnostic pretrained WM, LPS-mix* has significantly improved the base policy (BC) performance across tasks. Although HPT was pretrained on a large-scale dataset with more than 20 embodiments, such an embodiment-dependent pretrained policy performs poorly given a small amount of target-embodiment data for finetuning.
TABLE II: Latent Policy Steering in Robomimic
@@ V-B Real-world experiments
Transport 22.8±2.1 22.5±4.1 23.5±0.7 19.8±1.2 23.2±0.5 25.8±1.6 30.4±3.1 27.9±6.2 32.9±2.8 34.6±3.6
Average 33.2% 34.0% 35.9% 34.3% 35.5% 57.3% 52.9% 62.4% 62.9% 63.4%
We report the mean success rate across 3 seeds with standard deviations. Despite the majority of the pretrain data are collected in the real-world, our proposed method (e.g., LPS-mix*) achieves 10.6% relative improvement based on the behavior cloned policy in the simulation, 50 demonstrations, averaged across 4 tasks. Since transport is a long-horizon bimanual task with a greater action dimension and requiring precision, the LPS’s significant improvement over the baseline (34% relative improvement) with sufficient demonstrations shows its potential beyond simple tasks and single-arm robot systems. The improvement from LPS with or without pretraining is less significant, given 30 demonstrations. We believe LPS performs best when given a multimodal policy (i.e., a policy that generates diverse plans), whereas 30 demonstrations could lead to unimodal policies that are difficult to improve.
Does the embodiment-agnostic pretrained WM leverage multi-embodiment data more effectively than an embodiment-dependent pretrained policy? We evaluate LPS-mix* against three other baselines (BC, LPS*, HPT) in the real world on a pick-and-place task and three challenging tasks involving tool-use, and deformable object manipulation. LPS-mix* significantly improves the policy (BC) performance across tasks (Table I) (relative improvement of 70% and 44%, given 30-50 and 60-100 demonstrations). While HPT has been pretrained on a large-scale dataset with more than 20 embodiments, its embodiment-dependent pretraining makes it challenging to finetune to a target-embodiment in a low-data regime. On the other hand, by using an embodiment-agnostic pretrained WM, LPS can translate an extra multi-embodiment dataset into better performance on an unseen target embodiment during pretraining, with only a small amount of data. Both BC and LPS* that learn from scratch perform poorly, possibly due to a lack of data.
V-C Ablation studies in the simulated benchmark
@@
We compare the WM pretrained with embodiment-agnostic actions (optical flow) against the WM pretrained with end-effector action (eef). The optical-flow-based WM produces better results on a dataset of 3 robots. With more embodiments available, we believe the performance gap will widen.
Figure 3: Scaling through embodiments. We evaluate LPS in the Robomimic given different pretrain mixtures. LPS-mix manages to improve the behavior cloning policies’ performance up to 10.6%, given 50 target-embodiment demonstrations, averaged across four tasks, considering simulation, real-world robot data, and human videos from play.
How does LPS’s performance vary with different embodiments used during pretraining? Given four Robomimic tasks, we compare LPS-mix* against LPS variants pretrained with real-world robot data (LPS-real*) or simulation data (LPS-sim*), and baselines without pretraining (BC, LPS*), given 30-50 demonstrations. In simulations, we expect the performance improvement due to the embodiment-agnostic pretrained WM to be lower than in real-world experiments, given that the pretrain mixture of LPS-real* and LPS-mix* contain primarily real-world data.
Given 50 demonstrations, LPS-mix* outperforms other LPS variants and improves a behavior cloned policy by 10.6%, despite the majority of its pretrain data being collected in the real world (robot and human embodiments). LPS’s results regarding the transport task are particularly encouraging, since it is a long-horizon bimanual task with a higher action dimension (14) that requires coordination between two arms during transport. It demonstrates LPS’s potential beyond simple tasks and single-arm robot systems.
LPS’s improvement over the BC baseline is less significant in the 30 demonstration cases, regardless of pretraining mixture. In our opinion, LPS works best when given a multimodal policy that generates diverse plans. Too few demonstrations will result in a unimodal policy that generates similar plans and gives little room for improvement.
How does LPS’s performance scale with the number of embodiments? We explore how different combinations/numbers of embodiments for pretraining affect the performance, averaged across four Robomimic tasks, as shown in Fig. 3. Our target dataset has 50 demonstrations, and the pretraining dataset could include simulation robot data, real-world robot data, and human video data from play. We observe a promising trend: thanks to the embodiment-agnostic action representation and WM, increasing the number of embodiments leads to more usable pretraining data and results in better performance compared to the behavior-cloned baseline. While embodiment-dependent pretrained policies such as HPT [35] could also scale with embodiments, we do not compare to it since HPT pretrained with more than 20 embodiments performs sub-optimally given a small amount of finetuning data on the target embodiment, as shown in real-world experiments (Table I).
Surprisingly, we observe that the human video data from play leads to a very competitive pretrained WM, despite having less data compared to other data mixtures with more embodiments. Our hypothesis is that humans are able to demonstrate more diverse and cohesive manipulation skills with their hands in a shorter period of time than when teleoperating a robot, resulting in higher data quality.
@@
How does the optical flow compare to other off-the-shelf action representations? Robot end-effector pose (EEF) is another off-the-shelf action representation for manipulators that can be moderately embodiment-agnostic. In simulation, we investigate the advantage of using optical flow as an off-the-shelf embodiment-agnostic action representation during pretraining over EEF pose, given a pretraining dataset with 3 different robots. Our approach (LPS-sim(flow)*) uses optical flow action during pretraining, and LPS-sim(eef) uses EEF action during pretraining. Both are finetuned on the target dataset with 30 or 50 demonstrations, using the EEF robot action.
@@
The results are shown in Table III. The advantage of optical flow as embodiment-agnostic actions vs. EEF actions is clear, given 3 embodiments. We suspect the advantage will keep growing when the dataset size is larger and the number of embodiments is larger. Notice that optical flow also allows human embodiment to be part of pretraining, while EEF actions are only available for robots.
TABLE IV: Latent Policy Steering’s value function ablations
@@
Transport 27.3±2.1 30.5±4.5 29.6±2.9 35.3±6.4
Average 62.9% 65.2% 64.3% 68.7%
We ablate our designs of the robust value function for LPS into 2 variants (LPS-vanilla/bootstrap). In both variants, performance is lower than the policy-only baseline (BC), demonstrating the necessity of designing a robust value function to let WM benefit the policy (BC).
How does the robust value function affect LPS’s performance? LPS uses a robust value function to steer the policy back to the expert dataset. Such a value function is trained on data from the expert dataset and the data likely to be visited by the policy during inference. We ablate these designs through 2 variants below.
  • LPS-vanilla: This variant only trains the value function on the data from the expert dataset, given environmental rewards. It will select the best candidate based on the value function.
  • LPS-bootstrap: This variant learns a value function both on the expert dataset and the data likely to be visited by the policy during inference. It achieves this by simulating the states likely visited by the policy through the WM. Different from the LPS, the extra states are labeled with the same rewards as the dataset’s states and do not include the reward penalizing distribution shifts, modifying line 8 of Alg. 1 ($r^{\prime}_{t:t+h}=r_{t:t+h}$). It will select the best candidate based on the value function.
All models use 100 demonstrations to train from scratch and no pretraining is involved.
Figure 4: Latent Policy Steering with different horizons. We found that the proposed LPS method can effectively improve policy performance from a shorter horizon to a longer horizon up to 16, both learned from scratch. When the horizon becomes very long (e.g., 20, 24), the reward used to capture distribution shift becomes noisy, making the value function less useful during inference.
How does the inference horizon affect LPS’s performance? We investigate LPS’s performance given different action prediction horizons from small to large. The WM’s prediction horizon always matches the action prediction horizon of the base diffusion policy [8]. We learn the policy and the WM from scratch, given EEF action, with 100 demonstrations on the Can task of Robomimic [22].
We found that LPS with a WM performs better than BC at horizons 4,8,16,20, and performs worse than BC for a horizon of 24 (Fig. 4). Large action prediction size/horizon of 24 leads to a suboptimal policy with poor performance, and the penalization on deviation from the dataset will be noisy (lines 8-10, Alg. 1). This results in a noisy reward and makes the value function less useful.
VI CONCLUSIONS & DISCUSSION
We pretrain an embodiment-agnostic World Model (WM) on existing or cost-effective data sources such as public multi-embodiment robot datasets [34], robot data from simulations [42], and easily collected human data from play to improve visuomotor policies with a small amount of real-world data. We propose optical flow as an embodiment-agnostic action representation to enable this training across diverse embodiments. The resulting pretrained WM has minimal dependency on specific embodiments and can be easily finetuned to a target embodiment with a small amount of data. Our proposed algorithm, Latent Policy Steering, manages to improve a policy’s performance in the real-world evaluations across 4 tasks relatively by 70% with 30-50 expert demonstrations, and 44% with 60-100 expert demonstrations, showing its effectiveness over prior work without pretraining a WM.
Despite the advantages of being embodiment-agnostic and easily computable, we observe several limitations with optical flow as an action representation. Firstly, optical flow cannot reliably capture motions if an occlusion occurs. Furthermore, optical flow depends on the viewpoint. The same skill will lead to different optical flow patterns given different viewpoints. However, this limitation can be overcome by leveraging data with diverse viewpoints, such as that present in large cross-embodiment datasets. In addition, visual observations from non-static cameras (e.g., mobile robots) can produce noisy optical flow without further processing. We plan to investigate how to build more complete and scalable action representations by bridging different embodiments with embodiment-agnostic action and augmenting it with embodiment-specific details when the previous action representation lacks critical information for actions.